# Heuristic Adaptation of Potentially Misspecified Domain Support for Likelihood-Free Inference in Stochastic Dynamical Systems

### v2

@@ III-D Why not use the widest support possible?
IV Heuristic support optimisation methods
@@ IV-B BayesSim-MODE: mode-shift-based expansion
17:     $\Delta_{L}^{(d)}\leftarrow[\theta_{\min}^{(d)},\theta_{\min}^{(d)}+\delta]$ $\triangleright$ Left edge zone
18:     $\Delta_{R}^{(d)}\leftarrow[\theta_{\max}^{(d)}+\delta,\theta_{\max}^{(d)}]$ $\triangleright$ Right edge zone
19:     $W_{L}^{(d)}\leftarrow 0$; $W_{R}^{(d)}\leftarrow 0$
20:     for each Gaussian $k\in$ MoG do
21:      $\nu_{i,k}^{(d)}\leftarrow\mu_{i,k}^{(d)}-\mu_{i-1,k}^{(d)}$ $\triangleright$ Mode shift
22:      // Normalised proximity to bounds
23:      $z_{k}^{(d)}\leftarrow\frac{(\mu_{i,k}^{(d)}-\theta_{\min}^{(d)})}{r^{(d)}}$
24:      // Check mode shift & bound proximity
@@
28:        $W_{R}^{(d)}\leftarrow W_{R}^{(d)}+w_{i,k}$
29:      end if
30:     end for
31:     if $W_{L}^{(d)}>\tau\And(\theta_{\min}^{(d)}-\eta r^{(d)})\in\Phi$ then
32:      $\theta_{\min}^{(d)}\leftarrow\theta_{\min}^{(d)}-\eta r^{(d)}$ $\triangleright$ Expand left
33:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\min}^{(d)}\leftarrow\Phi_{\min}^{(d)}$
34:     if $W_{R}^{(d)}>\tau\And(\theta_{\max}^{(d)}+\eta r^{(d)})\in\Phi$ then
35:      $\theta_{\max}^{(d)}\leftarrow\theta_{\max}^{(d)}+\eta r^{(d)}$ $\triangleright$ Expand right
36:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\max}^{(d)}\leftarrow\Phi_{\max}^{(d)}$
37:     $d\leftarrow d+1$
38:   end while
39:   $i\leftarrow i+1$
40:   $\Theta_{i}\leftarrow[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$ $\triangleright$ Update bounds
41:   $p_{i}\leftarrow\hat{p}_{(\Theta_{i})}$ $\triangleright$ Update reference prior
42:  end while
43:  return $\hat{p}$
44:end function
@@
Similarly to EDGE, the standard BayesSim assumptions apply. We also have MODE’s hyperparameters, which are the mode shift threshold $\nu_{\text{TH}}$, the bounds proximity threshold $\rho$, the MoG weights summary threshold $\tau$, the bounds expansion factor $\eta$, and the physically feasible domain limits $\Phi$. The main difference between MODE and EDGE is that we now calculate the shift of the MoG posterior modes $\boldsymbol{\mu}_{i}$ compared to the previous iteration $\boldsymbol{\mu}_{i-1}$ (line 21) along with their normalised proximity to the bounds (line 23). Thus, we do not check for $\Theta$ adaptation at the end of the first inference iteration. In all subsequent iterations, the MoG details of the previous iteration are accessible through the reference prior.
We accumulate the mixture coefficients (weights) of each mode that satisfies the shift and proximity thresholds for the left or right bound (lines 26, LABEL: and 28) of a parameter $d$. The accumulated weights $W_{L}^{(d)}$, $W_{R}^{(d)}$ are used for the expansion criteria (lines 31, LABEL: and 34). If $\tau$ has been exceeded and if an expansion would be within $\Phi$ (lines 31, LABEL: and 34), we expand it by a factor of $\eta$ (lines 32, LABEL: and 35). If $\tau$ has been exceeded, but the expansion would have been beyond the corresponding $\Phi$ limit, then this becomes the new boundary.
@@ V-B Results and discussion
For Lotka-Volterra (fig. 7), we see that with a good support (OK), the inference result consistently approximates the ground-truth $X,Y$ population trajectories. CENTRE and EDGE perform consistently better than using broad support. Although CENTRE is better in terms of trajectory proximity to ground truth, the per-parameter plots show that EDGE is more robust in per-parameter prediction accuracy. MODE underperforms, with its temporal shift heuristic being difficult to tune, especially given the task’s stochastic dynamics. Thus, it is almost as inaccurate as the misspecified support inference. These results also motivate a more thorough exploration of how each physical parameter’s value influences the resulting population trajectories. Qualitatively, our results suggest that $\boldsymbol{\theta}_{2}$ (a predator dies) and $\boldsymbol{\theta}_{4}$ (a prey dies) inference performance is better correlated with ground truth trajectory approximation, followed by $\boldsymbol{\theta}_{3}$ (a prey is born).
For M/G/1 (fig. 8), it is qualitatively evident that its increased stochasticity poses a bigger challenge for LFI. Although MODE performs better overall in per-parameter inference, EDGE is more consistent among iterations, which is evident in the proximity of the calibrated trajectory to ground truth. We also see the adversarial effect of a broad support mainly for $\boldsymbol{\theta}_{1}$ and $\boldsymbol{\theta}_{2}$ inference. $\boldsymbol{\theta}_{3}$ seems to be inferred more robustly overall, possibly because in the chosen domain $\boldsymbol{\theta}_{3}$ perturbations have a smaller impact on system trajectory. We also see that among all the M/G/1 $\boldsymbol{\theta}$ components, the $\boldsymbol{\theta}_{1}$ inference generally has a stronger qualitative correlation with the overall trajectory approximation.
@@ VI Real2Sim2Real for visuomotor DLO whipping
Following our inference results for Lotka-Volterra and M/G/1, we use BayesSim-EDGE to study a higher-dim dynamical system. This is the prefaced visuomotor DLO whipping task with a sparse outcome-dependent reward that implicitly guides the entire DLO body toward a stack of cubes. We use the integrated Real2Sim2Real framework for visuomotor DLO manipulation presented in [46] (fig. 9).
@@ VII-C Real2Sim2Real for object-centric agent adaptation
The overall scores give further insight into the impact of support adaptation on object-centric agent performance. In the default support MoG posteriors of fig. 12, we see that DLO-0 needs an adaptation of at least the stiffness subdomain. BayesSim-EDGE achieves this, along with some proportionately smaller length subdomain adaptation, leading to increased performance for the corresponding DLO-$0^{\prime}$ policy. In contrast, DLO-$1$ does not need any support adaptation, but BayesSim-EDGE results in a sharper posterior, leading to an approximately median performance for PPO-$1^{\prime}$. The MoG inferred by BayesSim-EDGE for DLO-$2$ leads to PPO-$2^{\prime}$, which overall is more useful than PPO-$2$. In the DLO-$3$ EDGE posterior, we see a length adaptation. However, the latent interaction of the length subdomain adaptation with the Young’s and density subdomains results in PPO-$3^{\prime}$ being even more useful for DLO-$2$ than DLO-$3$.

### v3

@@ III-D Why not use the widest support possible?
IV Heuristic support adaptation methods
@@ IV-B BayesSim-MODE: mode-shift-based expansion
17:     $W_{L}^{(d)}\leftarrow 0$; $W_{R}^{(d)}\leftarrow 0$
18:     for each Gaussian $k\in$ MoG do
19:      $\nu_{i,k}^{(d)}\leftarrow\mu_{i,k}^{(d)}-\mu_{i-1,k}^{(d)}$ $\triangleright$ Mode shift
20:      // Normalised proximity to bounds
21:      $z_{k}^{(d)}\leftarrow\frac{(\mu_{i,k}^{(d)}-\theta_{\min}^{(d)})}{r^{(d)}}$
22:      // Check mode shift & bound proximity
@@
26:        $W_{R}^{(d)}\leftarrow W_{R}^{(d)}+w_{i,k}$
27:      end if
28:     end for
29:     if $W_{L}^{(d)}>\tau\And(\theta_{\min}^{(d)}-\eta r^{(d)})\in\Phi$ then
30:      $\theta_{\min}^{(d)}\leftarrow\theta_{\min}^{(d)}-\eta r^{(d)}$ $\triangleright$ Expand left
31:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\min}^{(d)}\leftarrow\Phi_{\min}^{(d)}$
32:     if $W_{R}^{(d)}>\tau\And(\theta_{\max}^{(d)}+\eta r^{(d)})\in\Phi$ then
33:      $\theta_{\max}^{(d)}\leftarrow\theta_{\max}^{(d)}+\eta r^{(d)}$ $\triangleright$ Expand right
34:     end $\triangleright if$ $\notin\Phi$ then $\theta_{\max}^{(d)}\leftarrow\Phi_{\max}^{(d)}$
35:     $d\leftarrow d+1$
36:   end while
37:   $i\leftarrow i+1$
38:   $\Theta_{i}\leftarrow[\boldsymbol{\theta}_{\min},\boldsymbol{\theta}_{\max}]$ $\triangleright$ Update bounds
39:   $p_{i}\leftarrow\hat{p}_{(\Theta_{i})}$ $\triangleright$ Update reference prior
40:  end while
41:  return $\hat{p}$
42:end function
@@
Similarly to EDGE, the standard BayesSim assumptions apply. We also have MODE’s hyperparameters, which are the mode shift threshold $\nu_{\text{TH}}$, the bounds proximity threshold $\rho$, the MoG weights summary threshold $\tau$, the bounds expansion factor $\eta$, and the physically feasible domain limits $\Phi$. The main difference between MODE and EDGE is that we now calculate the shift of the MoG posterior modes $\boldsymbol{\mu}_{i}$ compared to the previous iteration $\boldsymbol{\mu}_{i-1}$ (line 19) along with their normalised proximity to the bounds (line 21). Thus, we do not check for $\Theta$ adaptation at the end of the first inference iteration. In all subsequent iterations, the MoG details of the previous iteration are accessible through the reference prior.
We accumulate the mixture coefficients (weights) of each mode that satisfies the shift and proximity thresholds for the left or right bound (lines 24, LABEL: and 26) of a parameter $d$. The accumulated weights $W_{L}^{(d)}$, $W_{R}^{(d)}$ are used for the expansion criteria (lines 29, LABEL: and 32). If $\tau$ has been exceeded and if an expansion would be within $\Phi$ (lines 29, LABEL: and 32), we expand it by a factor of $\eta$ (lines 30, LABEL: and 33). If $\tau$ has been exceeded, but the expansion would have been beyond the corresponding $\Phi$ limit, then this becomes the new boundary.
@@ V-B Results and discussion
For Lotka-Volterra (fig. 7), we see that with a good support (OK), the inference result consistently approximates the ground-truth $X,Y$ population trajectories. CENTRE and EDGE perform consistently better than using broad support. Although CENTRE is better in terms of trajectory proximity to ground truth, the per-parameter plots show that EDGE is more robust in per-parameter prediction accuracy. MODE underperforms, with its temporal shift heuristic being difficult to tune, especially given the task’s stochastic dynamics. Thus, it is almost as inaccurate as the misspecified support inference. These results also motivate a more thorough exploration of how each physical parameter’s value influences the resulting population trajectories. Qualitatively, our results suggest that $\theta_{2}$ (a predator dies) and $\theta_{4}$ (a prey dies) inference performance is better correlated with ground truth trajectory approximation, followed by $\theta_{3}$ (a prey is born).
For M/G/1 (fig. 8), it is qualitatively evident that its increased stochasticity poses a bigger challenge for LFI. Although MODE performs better overall in per-parameter inference, EDGE is more consistent among iterations, which is evident in the proximity of the calibrated trajectory to ground truth. We also see the adversarial effect of a broad support mainly for $\theta_{1}$ and $\theta_{2}$ inference. $\theta_{3}$ seems to be inferred more robustly overall, possibly because in the chosen domain $\theta_{3}$ perturbations have a smaller impact on system trajectory. We also see that among all the M/G/1 $\boldsymbol{\theta}$ components, the $\theta_{1}$ inference generally has a stronger qualitative correlation with the overall trajectory approximation.
@@ VI Real2Sim2Real for visuomotor DLO whipping
Following our inference results for Lotka-Volterra and M/G/1, we use BayesSim-EDGE to study a higher-dim dynamical system. This is the prefaced visuomotor DLO whipping task with a sparse outcome-dependent reward that implicitly guides the entire DLO body [54] toward a stack of cubes. We use the integrated Real2Sim2Real framework for visuomotor DLO manipulation presented in [46] (fig. 9).
@@ VII-C Real2Sim2Real for object-centric agent adaptation
The overall scores give further insight into the impact of support adaptation on object-centric agent performance. In the default support MoG posteriors of fig. 12, we see that DLO-0 needs an adaptation of at least the stiffness subdomain. BayesSim-EDGE achieves this, along with some proportionately smaller length subdomain adaptation, leading to increased performance for the corresponding PPO-$0^{\prime}$ policy. In contrast, DLO-$1$ does not need any support adaptation, but BayesSim-EDGE results in a sharper posterior, leading to an approximately median performance for PPO-$1^{\prime}$. The MoG inferred by BayesSim-EDGE for DLO-$2$ leads to PPO-$2^{\prime}$, which overall is more useful than PPO-$2$. In the DLO-$3$ EDGE posterior, we see a length adaptation. However, the latent interaction of the length subdomain adaptation with the Young’s and density subdomains results in PPO-$3^{\prime}$ being even more useful for DLO-$2$ than DLO-$3$.